Sasha owns two investments, A and B, that have a combined total value of 36,300 dollars. Investment A is expected to pay 22,000 dollars in 1 year(s) from today and has an expected return of 3.38 percent per year. Investment B is expected to pay 25,497 dollars in T years from today and has an expected return of 5.32 percent per year. What is T, the number of years from today that investment B is expected to pay 25,497 dollars? Round your answer to 2 decimal places (for example, 2.89, 14.70, or 6.00).
Answer #1

Total Value = Investment A + Investment B

36,300 = 22,000/(1.0338) + 25,497/(1.0532)T

15,019.29 = 25,497/(1.0532)T

Using log,

T = log(25,497/15,019.29)/log(1.0532)

T = 10.21 years